# Neutron-nucleus dynamics simulations for quantum computers

### 22 Feb 2024 v1

@@ Appendix A Definitions and Lemmas
where the right-hand side is understood to be the binary representation of $i$. For example, $O_{4}=|100\rangle\!\langle 100|$. Furthermore, define $\tilde{O}_{j}$ to be the Pauli string composed of $I$ and $Z$ operators such that the bits of $j$ determine if the operator at each position is $Z$ or $I$. For a concrete example, consider that for $j=6_{10}=110_{2}$, each $1$ is represented by $Z$, and each 00 by $I$. The operator is then given by
@@
[…]
Since $k=2$, each of the above operators can be thought of as two-qubit flips in a particular subspace. For example, $|000\rangle\!\langle 011|+|011\rangle\!\langle 000|$ acts like $IXX$ in the subspace spanned by $\{|000\rangle,|011\rangle\}$ since its action on any state in this subspace is the same as the action of $IXX$. This is a unifying feature—any $\ell$ operator can be thought of as a product of $I$ (on the unchanged qubits) and $X$ (on the $\ell$ flipped qubits) acting on a particular subspace. Thus, an alternate way to represent the distance operators is as a composition of a projection onto the subspace of interest, followed by a string consisting of $X$ and $I$. For example,
@@
A basis encoding can alternatively be represented using a sequence of flipped bits. This alternate representation is defined as $S_{n}$. Thus, for an encoding of size $2^{n}$, the alternate representation is of size $12^{n}-$ 1.
@@
Basis
Gray
Binary
$|0\rangle\leftrightarrow|1\rangle$
$1$
$1$
$|1\rangle\leftrightarrow|2\rangle$
$2$
$3$
$|2\rangle\leftrightarrow|3\rangle$
$1$
$1$
$|3\rangle\leftrightarrow|4\rangle$
$3$
$7$
$|4\rangle\leftrightarrow|5\rangle$
$1$
$1$
$|5\rangle\leftrightarrow|6\rangle$
$2$
$3$
$|6\rangle\leftrightarrow|7\rangle$
$1$
$1$
@@
Thus, the reversed code has the same structure as the existing code with the first half entries ending with $1$, and the second half ending with 00. The entries of alternate representation indicate which bits are flipped and therefore, is unaffected if $0\leftrightarrow 1$. Thus, the alternate representation for the reversed Gray code is the same as the original.
@@
Similarly, for the binary code, we act with a bit-wise addition modulo 2 between the binary representation of every element in the subsequence. The resulting binary string is then translated into a Pauli string of $\{I,X\}$ – each $1$ is mapped to $X$, and each 00 is mapped to $I$. For example,
@@
The alternate represenatation for the Gray and binary code on $n$ qubits can be expressed in the form
@@
where $P_{i}$ stands for $i$ and $12^{i}-$ 1 in the Gray and binary codes, respectively, and we refer to each $P_{i}$ as a pivot. The term pivot refers to the fact that about $P_{n}$, the alternate representation $S_{n}$ is symmetric.
@@
The last entry of $\mathcal{B}_{n-1}$ is $1^{n-1}$, and the first entry of $\mathcal{B}_{n-1}$ is $0^{n-1}$. In the first and second half of the overall code, the first bit is never flipped, and between the halves all bits are flipped. Since the decimal representation of $1^{n}$ is $12^{n}-$ 1, the alternate representation is then given by
@@
(b) Case $4$: The vertical wave section of the subsequence can be canceled out. We construct a new subsequence with the dotted section, $P_{n+1}$, and the striped section, which reduces to Case $3$.
@@
Figure 18: Proof of Lemma A.8 can be broken down into multiple cases. We label sections of the subsequence with different colors. $P_{i}$ stands for $i$ and $12^{i}-$ 1 in the Gray and binary codes respectively.
@@
where $P_{i}$ stands for $i$ and $12^{i}-$ 1 in the Gray and binary codes, respectively, and we refer to each $P_{i}$ as a pivot. The pivots $P_{n}$, $P_{n+1}$, and $P_{n}$ occur at indices $2^{n-1}$, $2^{n}$, and $3\cdot 2^{n-1}$, respectively. Consider an arbitrary subsequence
@@
[…]
@@
[…]
@@
[…]
@@
  • Case $1$: The shorter subsequence ends before index $2^{m}$. More concretely, the shorter subsequence exists within $[0,2^{m})$ and using Lemma A.8, there exists an equivalent subsequence that ends at index $2^{m-1}$. No possible subsequence in this region can have its corresponding Pauli string with $X$ acting on qubit $m+1$.
@@
$|H(N,K)|=\begin{cases}2^{n}+n2^{n-1}&K=0\\ 2^{n}+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&1\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1},\end{cases}$ (124)
@@
$k$
$1$
$2$
$4$
$8$
$1$
XIII
IXII
IIXI
IIIX
$2$
XXII
XIXI
XIIX
$3$
XXXI
XXIX
$4$
IXXI
IXIX
$5$
IXXX
$6$
XXXX
$7$
XIXX
$8$
IIXX

### 15 Mar 2026 v3

@@ Appendix A Definitions and Lemmas
where the right-hand side is understood to be the binary representation of $i$. For example, $O_{4}=|100\rangle\!\langle 100|$. Furthermore, define $\tilde{O}_{j}$ to be the Pauli string composed of $I$ and $Z$ operators such that the bits of $j$ determine if the operator at each position is $Z$ or $I$. For a concrete example, consider that for $j=6_{10}=110_{2}$, each $1$ is represented by $Z$, and each 0 by $I$. The operator is then given by
@@
$\{2,3\}$ $=\bigg\{|000\rangle\!\langle 011|+|011\rangle\!\langle 000|,|001\rangle\!\langle 010|+|010\rangle\!\langle 001|,$
$|100\rangle\!\langle 111|+|111\rangle\!\langle 100|,|101\rangle\!\langle 110|+|110\rangle\!\langle 101|.\bigg\},$
$\{1,2\}$ $=\bigg\{|000\rangle\!\langle 110|+|110\rangle\!\langle 000|,|010\rangle\!\langle 100|+|100\rangle\!\langle 010|,$
$|001\rangle\!\langle 111|+|111\rangle\!\langle 001|,|011\rangle\!\langle 101|+|101\rangle\!\langle 011|.\bigg\},$
$\{1,3\}$ $=\bigg\{|000\rangle\!\langle 101|+|101\rangle\!\langle 000|,|001\rangle\!\langle 100|+|100\rangle\!\langle 001|,$
$|010\rangle\!\langle 111|+|111\rangle\!\langle 010|,|011\rangle\!\langle 110|+|110\rangle\!\langle 011|.\bigg\}.$ (96)
Since $k=2$, each of the above operators can be thought of as two-qubit flips in a particular subspace. For example, $|000\rangle\!\langle 011|+|011\rangle\!\langle 000|$ acts like $IXX$ in the subspace spanned by $\{|000\rangle,|011\rangle\}$ since its action on any state in this subspace is the same as the action of $IXX$. This is a unifying feature—any distance-$\ell$ operator can be thought of as a product of $I$ (on the unchanged qubits) and $X$ (on the $\ell$ flipped qubits) acting on a particular subspace. Thus, an alternate way to represent the distance operators is as a composition of a projection onto the subspace of interest, followed by a string consisting of $X$ and $I$. For example,
@@
A basis encoding can alternatively be represented using a sequence of flipped bits. This alternate representation is defined as $S_{n}$. Thus, for an encoding of size $2^{n}$, the alternate representation is of size $2^{n}-1$.
@@
Basis Gray Binary
$|0\rangle\leftrightarrow|1\rangle$ 11 11
$|1\rangle\leftrightarrow|2\rangle$ 22 33
$|2\rangle\leftrightarrow|3\rangle$ 11 11
$|3\rangle\leftrightarrow|4\rangle$ 33 77
$|4\rangle\leftrightarrow|5\rangle$ 11 11
$|5\rangle\leftrightarrow|6\rangle$ 22 33
$|6\rangle\leftrightarrow|7\rangle$ 11 11
@@
Thus, the reversed code has the same structure as the existing code with the first half entries ending with $1$, and the second half ending with 0. The entries of alternate representation indicate which bits are flipped and therefore, is unaffected if $0\leftrightarrow 1$. Thus, the alternate representation for the reversed Gray code is the same as the original.
@@
Similarly, for the binary code, we act with a bit-wise addition modulo 2 between the binary representation of every element in the subsequence. The resulting binary string is then translated into a Pauli string of $\{I,X\}$ – each $1$ is mapped to $X$, and each 0 is mapped to $I$. For example,
@@
The alternate representation for the Gray and binary code on $n$ qubits can be expressed in the form
@@
where $P_{i}$ stands for $i$ and $2^{i}-1$ in the Gray and binary codes, respectively, and we refer to each $P_{i}$ as a pivot. The term pivot refers to the fact that about $P_{n}$, the alternate representation $S_{n}$ is symmetric.
@@
The last entry of $\mathcal{B}_{n-1}$ is $1^{n-1}$, and the first entry of $\mathcal{B}_{n-1}$ is $0^{n-1}$. In the first and second half of the overall code, the first bit is never flipped, and between the halves all bits are flipped. Since the decimal representation of $1^{n}$ is $2^{n}-1$, the alternate representation is then given by
@@
(b) Case $4$: The vertical wave section of the subsequence can be canceled out. We construct a new subsequence with the dotted section, $P_{n+1}$, and the striped section, which reduces to Case 33.
@@
Figure 20: Proof of Lemma A.8 can be broken down into multiple cases. We label sections of the subsequence with different colors. $P_{i}$ stands for $i$ and $2^{i}-1$ in the Gray and binary codes respectively.
@@
where $P_{i}$ stands for $i$ and $2^{i}-1$ in the Gray and binary codes, respectively, and we refer to each $P_{i}$ as a pivot. The pivots $P_{n}$, $P_{n+1}$, and $P_{n}$ occur at indices $2^{n-1}$, $2^{n}$, and $3\cdot 2^{n-1}$, respectively. Consider an arbitrary subsequence
@@
    In essence, we have replaced $P_{n+1}$ with $P_{n}$ in the original subsequence. This subsequence is now guaranteed to exist in $S_{n}=(S_{n-1},P_{n},S_{n-1})$. By the inductive hypothesis, there exists a subsequence $(a[k],\ldots,a[2^{n-1}-1],P_{n})$ of length $\leq\text{len}$, where $k\in[0,2^{n-1})$. By symmetry, the following subsequence
@@
    where we have replaced $P_{n}$ with $P_{n+1}$ again, must exist, has length $\leq\text{len}$, and ends at $n+1$.
@@
    The length of this subsequence is $\leq\text{len}$ and it ends at a power of two.
@@
  • Case $1$: The shorter subsequence ends before index $2^{m}$. More concretely, the shorter subsequence exists within $0,2^{m})$ and using Lemma A.8, there exists an equivalent subsequence that ends at index $2^{m-1}$. No possible subsequence in this region can have its corresponding Pauli string with $X$ acting on qubit $m+1$.
@@
Let $i,j$ be two $n$-bit binary strings, and let $N=2^{n}$. Then, the following statement is true
$P(n,K)\colon\sum_{i=0}^{2^{n}-1}(-1)^{i\cdot j}i^{K}=0\iff|j|>K,$ (125)
for all $0\leq K\leq N$ and $n>0$.
Proof. We prove the statement using induction. More concretely, we use simple induction on $n$ and strong induction on $K$. Thus, we assume $P(n-1,K),P(n-1,K-1),\ldots,P(n-1,0)$ to be true, and use them to prove $P(n,K)$. We first show the sets of base cases $P(1,K)$ and $P(n,0)$.
Base Case 1 : $P(1,K)$. The left hand side now becomes
$(-1)^{0\cdot j}0^{K}+(-1)^{1\cdot j}1^{K}$
$=0^{K}+(-1)^{j}=0\iff K=0,j=1,$ (126)
where we use the fact that $0^{0}=1$. Thus, $|j|>K$.
Base Case 2 : $P(n,0)$. The left hand side now becomes
$\sum_{i=0}^{2^{n}-1}(-1)^{i\cdot j}i^{0}=0,$ (127)
which is true $\forall j\neq 0$. Thus, $|j|>0$.
To prove the equivalence $P(n,K)$, we start from the right hand side $|j|>K$, and show that it is equivalent to the left hand side. Let the binary expansion of $j$ be given by $j_{1}j_{2}\ldots j_{n}$.
Case 1: $j_{1}=0$. Thus, $|j_{2}\ldots j_{n}|>K$. Using the inductive hypotheses and the fact that $|j_{2}\ldots j_{n}|>K$ satisfies the right hand side of all the hypotheses, we have the following equalities:
$\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K}$ $=0\qquad\text{Using }P(n-1,K)$ (128)
$\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-1}$ $=0\qquad\text{Using }P(n-1,K-1)$ (129)
[…]
$\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{0}$ $=0\qquad\text{Using }P(n-1,0).$ (131)
In each of the statements above, $i$ is an (n1)(n-1)-bit binary string. Now consider the following equation:
$\binom{K}{0}(2^{n})^{0}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K}\right]$
$+\binom{K}{1}2^{n}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-1}\right]$
$+\binom{K}{2}(2^{n})^{2}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-2}\right]$
$+\cdots+\binom{K}{K}(2^{n})^{K}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{0}\right]$
$=0.$ (132)
The equality is because each term within square brackets is zero (using (128)-(131)). Thus, using the binomial theorem,
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}(2^{n}+i)^{K}\right]=0.$ (133)
Combining with the first step of the inductive tree (128),
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}(i)^{K}\right]\\ +\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}(2^{n}+i)^{K}\right]=0.$ (134)
In the above two summations, $i$ is an n1n-1 bit binary string. We now create an $n$ bit binary string $i^{\prime}$ by appending either 0 or $1$ to the front of $i$ and append $j_{1}$ to the front of $j_{2}\ldots j_{n}$. Since $j_{1}=0$, this extra bit has no effect. Thus, the equation becomes
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{0i\cdot(j_{1}j_{2}\ldots j_{n})}(i)^{K}\right]\\ +\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{1i\cdot(j_{1}j_{2}\ldots j_{n})}(2^{n}+i)^{K}\right]=0.$ (135)
Now, we notice that $1i$ is the binary expansion of $2^{n}+i$. Thus, the summation can be written as
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{0i\cdot(j_{1}j_{2}\ldots j_{n})}(i)^{K}\right]\\ +\left[\sum_{i^{\prime}=2^{n-1}}^{2^{n}-1}(-1)^{i^{\prime}\cdot(j_{1}j_{2}\ldots j_{n})}(i^{\prime})^{K}\right]=0.$ (136)
Since $i^{\prime}$ is just a dummy index, we replace it with $i$ and combine it with the first term, finally leading to
$\sum_{i=0}^{2^{n}-1}(-1)^{i\cdot j}i^{K}=0.$ (137)
Case 2: $j_{1}=1$. Thus, $|j_{2}\ldots j_{n}|>K-1$. Using the inductive hypotheses and the fact that $|j_{2}\ldots j_{n}|>K-1$ satisfies the right hand side of all the hypotheses, we have the following equalities:
$\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-1}$ $=0\qquad\text{Using }P(n-1,K-1)$ (138)
$\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-2}$ $=0\qquad\text{Using }P(n-1,K-2)$ (139)
[…]
$\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{0}$ $=0\qquad\text{Using }P(n-1,0).$ (141)
In each of the statements above, $i$ is an (n1)(n-1)-bit binary string. Now consider the following equation:
$-\binom{K}{1}2^{n}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-1}\right]$
$-\binom{K}{2}(2^{n})^{2}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-2}\right]$
$-\cdots-\binom{K}{K}(2^{n})^{K}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{0}\right]$
$=0.$ (142)
The equality is because each term within square brackets is zero (using (138)-(141)). Adding and subtracting the following term,
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K}\right],$ (143)
we get
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K}\right]$
$-\left\{\binom{K}{0}(2^{n})^{0}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K}\right]\right$.
$+\binom{K}{1}2^{n}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-1}\right]$
$+\binom{K}{2}(2^{n})^{2}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K-2}\right]$
$+\left.\cdots+\binom{K}{K}(2^{n})^{K}\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{0}\right]\right\}=0.$ (144)
Combining the terms within the curly brackets using the binomial theorem, we get
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}i^{K}\right]\\ -\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{i\cdot(j_{2}\ldots j_{n})}(2^{n}+i)^{K}\right]=0.$ (145)
In the above two summations, $i$ is an n1n-1 bit binary string. We now create an $n$ bit binary string $i^{\prime}$ by appending either 0 or $1$ to the front of $i$ and append $j_{1}$ to the front of $j_{2}\ldots j_{n}$. Since $j_{1}=1$, this extra bit accounts for the negative sign. Thus, the equation becomes
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{0i\cdot(j_{1}j_{2}\ldots j_{n})}(i)^{K}\right]\\ +\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{1i\cdot(j_{1}j_{2}\ldots j_{n})}(2^{n}+i)^{K}\right]=0.$ (146)
Now, we notice that $1i$ is the binary expansion of $2^{n}+i$. Thus, the summation can be written as
$\left[\sum_{i=0}^{2^{n-1}-1}(-1)^{0i\cdot(j_{1}j_{2}\ldots j_{n})}(i)^{K}\right]\\ +\left[\sum_{i^{\prime}=2^{n-1}}^{2^{n}-1}(-1)^{i^{\prime}\cdot(j_{1}j_{2}\ldots j_{n})}(i^{\prime})^{K}\right].$ (147)
Since $i^{\prime}$ is just a dummy index, we replace it with $i$ and combine it with the first term, finally leading to
$\sum_{i=0}^{2^{n}-1}(-1)^{i\cdot j}i^{K}=0.$ (148)
Thus, for both cases $j_{1}=0,1$, we showed that $P(n,K)$ follows from the inductive hypotheses, concluding the proof.
Lemma A.11
Consider an $N\times N$ matrix $H$ with diagonal entries of the form
$H_{i,i}=\sum_{k=0}^{K}a_{k}i^{k},$ (149)
for some set $\{a_{k}\}_{k=0}^{K}$ of real constants and for all $i\in\left\{0,\ldots,N-1\right\}$. Considering only the diagonal part of the matrix, and using the binary encoding, we get the following equality:
$H_{\operatorname{diag}}$ $=\sum\limits_{i=0}^{N-1}\sum_{k=0}^{K}a_{k}i^{k}|i\rangle\!\langle i|$
$=\sum\limits_{i=0}^{N-1}\sum_{k=0}^{K}a_{k}i^{k}O_{i}$
$=\frac{1}{2^{n}}\sum\limits_{j=0}^{N-1}\sum_{k=0}^{K}a_{k}\left[\sum\limits_{i=0}^{N-1}(-1)^{i\cdot j}i^{k}\right]\tilde{O}_{j},$ (150)
where the third equality follows from Lemma A.1. Then, the number of Pauli terms accounted for by the diagonal part of the matrix is given by
$\operatorname{d}(n,K)\coloneqq\sum_{m=0}^{K}\binom{n}{m}$, (151)
where $n=\log_{2}(N).$
Proof. In (150), using Lemma A.10, the term in square brackets is non-zero if and only if $|j|\leq K$ and all terms with $|j|>K$ get cancelled. Thus, the surviving terms have $|j|\leq K$ and the number of such terms is given by
$\sum_{m=0}^{K}\binom{n}{m}$, (152)
which is the number of ways to pick $j$ such that $|j|\leq K$.
Remark 4
Lemma A.11 shows that the number of Pauli terms for the diagonal part of the Hamiltonian, using the binary encoding, is given by $\operatorname{d}(n,K)$. Instead, if we use the Gray encoding, the diagonal part of the Hamiltonian is given by
$H_{\operatorname{diag}}$ $=\sum\limits_{i=0}^{N-1}\sum_{k=0}^{K}a_{k}i^{k}|i\rangle\!\langle i|$
$=\sum\limits_{i=0}^{N-1}\sum_{k=0}^{K}a_{k}i^{k}O_{i}$
$=\frac{1}{2^{n}}\sum\limits_{j=0}^{N-1}\sum_{k=0}^{K}a_{k}\left[\sum\limits_{i=0}^{N-1}(-1)^{e(i)\cdot j}i^{k}\right]\tilde{O}_{j}.$ (153)
Replacing $i$ with $e(i)$ in the exponent changes the relative signs of each term. Thus, the set of surviving $j$ is different for the binary and Gray encodings. However, since we consider a sum of all possible $j$, the size of the set of surviving terms is the same and is equal to $\operatorname{d}(n,K)$.
Lemma A.12
@@
$|H(N,K)|=\begin{cases}\operatorname{d}(n,1)+n2^{n-1}&K=0\\ \operatorname{d}(n,K)+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&1\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1},\end{cases}$ (154)
@@
$k$ 11 22 44 88
11 XIII IXII IIXI IIIX
22 XXII XIXI XIIX
33 XXXI XXIX
44 IXXI IXIX
55 IXXX
66 XXXX
77 XIXX
88 IIXX